In a certain​ country, the true probability of a baby being a girl is 0.473 . Among the next four randomly selected births in the​ country, what is the probability that at least one of them is a ​boy?
The probability is
1
Expert's answer
2020-11-11T20:00:18-0500

probability that at least one of them is a ​boy = 1 - (probability that all of them are girls) = 1 - 0.4734 = 0.9499